OpenStudy (anonymous):

what must be added to x3 - 3x2 - 12x + 19 so that the result is exactly divisible by x2 + x -6 ?

OpenStudy (akashdeepdeb):

WELCOME TO OPENSTUDY!! :D

OpenStudy (anonymous):

divide dem both

OpenStudy (anonymous):

Is it your question ?: 3x - 6 - 12x +19 ? If we have that can write it : -9x+9 And can write the 2th sentence this : 3x-6 Now we should reduce the 2th from 1th sentence . It s : 3x-6-(-9x+9)=3x-6+9x-9=12x-15 Understand?!:)

OpenStudy (anonymous):

@E.ali dude u had interpreted wrongly actually question is x^3-3x^2-12x+19 -------------equation 1 x^2+x-6 --------equation 2 @queen151999 Ans will be equation 1 / equation 2 which will be equal to -2x-5

OpenStudy (debbieg):

No, not quite. The question is: if f(x)= \(x^3-3x^2-12x+19 \) & g(x)=\(x^2+x-6 \) then what must be ADDED to f(x) so that the division f(x)/g(x) has NO remainder?

OpenStudy (debbieg):

since \(x^2+x-6 =(x+3)(x-2)\), those will both need to be factors of the revised f(x) when you're finished with it.

OpenStudy (debbieg):

I think the idea is just to find a factor (x-a) so that \((x+3)(x-2)(x-a)=x^3-3x^2-12x+19+h(x)\) The easiest way is just to pick an a, find that product, and then see what you need to add to \(x^3-3x^2-12x+19\) to GET that. Whatever you have to add, that will be your h(x).

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."